# Chow polynomials of totally nonnegative matrices and posets

### v2

@@ Abstract.
Huh-Stevens and Ferroni-Schröter independently conjectured that Hilbert-Poincaré series of Chow rings of geometric lattices have only real zeros. Ferroni, Matherne and the second author extended this conjecture to Chow polynomials of Cohen-Macaulay poset. In this paper we address the above conjectures by providing new defining relations and properties of Chow functions of posets and matrices. These are used, in conjunction with new techniques on interlacing sequences of polynomials, to prove that Chow polynomials of totally nonnegative matrices have only real zeros, which, in turn, proves the above conjectures for a class of posets that contains projective and affine geometries, face lattices of cubical polytopes, dual partition and Dowling lattices, perfect matroid designs, and lattices of flats of paving matroids.
@@ 1. Introduction
In this paper we develop a systematic approach to the study of zeros of Chow polynomials for a class of posets called $\operatorname{\mathrm{TN}}$-posets, which were introduced in [BL24, BL25]. This class includes (rank-selected subposets of) affine and projective geometries, perfect matroid designs, dual partition lattices and dual Dowling lattices. To do so, we define and study Chow polynomials of totally nonnegative matrices ($\operatorname{\mathrm{TN}}$-matrices), i.e., matrices whose minors are all nonnegative. Our main result is Theorem 4.18, which states that the Chow polynomial of any lower triangular $\operatorname{\mathrm{TN}}$-matrix with all diagonal entries equal to one is real-rooted. A direct consequence of this is Theorem 6.1, which says that the Chow polynomial of any $\operatorname{\mathrm{TN}}$-poset is real-rooted. We also use our methods to prove that the Chow polynomial of any paving matroid is real-rooted, Corollary 7.2.
@@
In Section 5 we provide an explicit interpretation of the coefficients of the $\gamma$-polynomial of Chow polynomials of matrices as sums of certain minors of the matrix (Corollary 5.2) and as enumerators of certain non-intersecting paths in a network, following the work by Gessel and Viennot [GV85]. In Sections 6 and 7 we prove Theorem 6.1 and Theorem 7.1, which assert that (augmented) Chow polynomials of $\operatorname{\mathrm{TN}}$-posets and generalized paving posets (a class of posets that contains lattices of flats of paving matroids) are real-rooted.
@@ 2. The incidence algebra and Chow functions
In this section we recall the construction of Chow functions and Chow polynomials of partially ordered sets (posets). For undefined poset terminology, we refer to [Sta12]. Recall that an interval of a poset is a subposet of $P$ of the form $x,y]=\{z\in P:z\leq z\leq y\}$, where $x,y\in P$. All posets $P$ considered in this paper are locally finite, i.e., each interval of $P$ is finite. We also say that a poset is bounded if it has unique least and largest elements, which we denote by $\hat{0}$ and $\hat{1}$, respectively. Given a poset $P$, a weak rank function is a function $\rho:P\times P\to\mathbb{N}$ such that
@@
The function $g$ is called the left Kazhdan-Lusztig-Stanley function associated to $\kappa$.
@@ 4.2. Interlacing zeros and $\mathcal{I}_{n}$-interlacing sequences of polynomials
A polynomial $f\in\mathbb{C}[t]$ is called stable if either $f\equiv 0$ or all the zeros of $f$ have nonpositive imaginary parts. The Hermite-Biehler theorem relates stability to interlacing zeros, for a proof see [Wag92] [^4] There is a typo in the definition of proper position in [Wag92]. It is essential to add the condition that the zeros of $f$ and $g$ interlace..
@@ 5.2. Interlacing properties of $\gamma$-Chow polynomials
6. Chow polynomials of $\mathrm{TN}$-posets
@@
    $q$-poset with nonnegative $h$-vectors [Ald10]. For example shellable $q$-complexes [GPR22].
@@
Suppose $P$ is a $\operatorname{\mathrm{TN}}$-poset of rank $n$. Then the characteristic Chow polynomial $H_{P}$ is real-rooted. Moreover $H_{\tau(P)}\prec H_{P}$ and $H_{[\hat{0},x]}\prec H_{P}$ for any $x$ of rank n1n-1.
@@
Let $R=R(P)$. By Proposition 3.5, $H_{n}=H_{P}$, $d_{n}=tH_{\tau(P)}$ and $H_{n-1}=H_{[\hat{0},x]}$. Hence the theorem follows from Theorem 4.18. ∎
@@
If we apply Theorem 6.1 to a above, then we recover the result of Hoster and Stump [HS25].
It follows from [BL25, Corollary 3.10] that $\mathrm{aug(P)}$ is $\operatorname{\mathrm{TN}}$ whenever $P$ is $\operatorname{\mathrm{TN}}$. Hence the above results are also true for the characteristic augmented Chow polynomials.
Theorem 6.3.
Suppose $P$ is a $\operatorname{\mathrm{TN}}$-poset of rank $n$. Then $G_{P}$ is real-rooted. Moreover $G_{\tau(P)}\prec G_{P}$ and $G_{[\hat{0},x]}\prec G_{P}$ for any $x$ of rank n1n-1.
Recall that the characteristic augmented Chow polynomial of a poset $P$ is equal to the one of its dual $P^{*}$ by [BV25, Corollary 4.2], from which the following consequence follows.
@@
Let $P$ be a $\operatorname{\mathrm{TN}}$-poset. Then the characteristic augmented Chow polynomial of its dual poset, $G_{P^{*}}$, is real-rooted. In particular if $L$ is a Dowling lattice, then $G_{L}$ is real-rooted.
@@ 8.1. Binomial and Sheffer posets
$\sum_{n=0}^{\infty}G_{n}(t)\frac{z^{n}}{C(n)}$ $=\frac{c(z)\cdot b(tz)-tb(tz)\cdot b(z)}{b(tz)-tb(z)}.$

### v3

@@ Abstract.
Huh-Stevens and Ferroni-Schröter independently conjectured that Hilbert-Poincaré series of Chow rings of geometric lattices have only real zeros. Ferroni, Matherne and the second author extended this conjecture to Chow polynomials of Cohen-Macaulay poset. In this paper we address the above conjectures by providing new defining relations and properties of Chow functions of posets and matrices. These are used, in conjunction with new techniques on interlacing sequences of polynomials, to prove that Chow polynomials of totally nonnegative matrices have only real zeros, which, in turn, proves the above conjectures for a class of posets that contains projective and affine geometries, face lattices of cubical polytopes, partition lattices and Dowling lattices, perfect matroid designs, and lattices of flats of paving matroids.
@@ 1. Introduction
In this paper we develop a systematic approach to the study of zeros of Chow polynomials for a class of posets called $\operatorname{\mathrm{TN}}$-posets, which were introduced in [BL24, BL25]. This class includes (rank-selected subposets of) affine and projective geometries, perfect matroid designs, dual partition lattices and dual Dowling lattices. To do so, we define and study Chow polynomials of totally nonnegative matrices ($\operatorname{\mathrm{TN}}$-matrices), i.e., matrices whose minors are all nonnegative. Our main result is Theorem 4.18, which states that the Chow polynomial of any lower triangular $\operatorname{\mathrm{TN}}$-matrix with all diagonal entries equal to one is real-rooted. A direct consequence of this is Theorem 6.1, which says that the Chow polynomial of any $\operatorname{\mathrm{TN}}$-poset (or dual of any $\operatorname{\mathrm{TN}}$-poset) is real-rooted. We also use our methods to prove that the Chow polynomial of any paving matroid is real-rooted, Corollary 7.2.
@@
In Section 5 we provide an explicit interpretation of the coefficients of the $\gamma$-polynomial of Chow polynomials of matrices as sums of certain minors of the matrix (Corollary 5.2) and as enumerators of certain non-intersecting paths in a network, following the work by Gessel and Viennot [GV85]. In Sections 6 and 7 we prove Theorem 6.1 and Theorem 7.1, which assert that (augmented) Chow polynomials of $\operatorname{\mathrm{TN}}$-posets, dual $\operatorname{\mathrm{TN}}$-posets and generalized paving posets (a class of posets that contains lattices of flats of paving matroids) are real-rooted.
@@ 2. The incidence algebra and Chow functions
In this section we recall the construction of Chow functions and Chow polynomials of partially ordered sets (posets). For undefined poset terminology, we refer to [Sta12]. Recall that an interval of a poset is a subposet of $P$ of the form $x,y]=\{z\in P:x\leq z\leq y\}$, where $x,y\in P$. All posets $P$ considered in this paper are locally finite, i.e., each interval of $P$ is finite. We also say that a poset is bounded if it has unique least and largest elements, which we denote by $\hat{0}$ and $\hat{1}$, respectively. Given a poset $P$, a weak rank function is a function $\rho:P\times P\to\mathbb{N}$ such that
@@
The function $g$ is called the left Kazhdan-Lusztig-Stanley (KLS) function associated to $\kappa$.
@@ 2.1. Augmented Chow functions
2.2. Duality for Chow functions
If $P$ is a finite weakly ranked poset, then so is its dual $P^{*}$, i.e., the poset on the same ground set as $P$ for which $x\leq_{P^{*}}y$ if and only if $y\leq_{P}x$. If $f\in I(P)$, let $f^{T}\in I(P^{*})$ be defined by $f^{T}_{y,x}=f_{x,y}$ for all $x\leq y$ in $P$. If $g$ is a left KLS function for $P$, then $g^{T}$ is a left KLS function for $P^{*}$. Denote by $H^{*},d^{*},G^{*}$ and $A^{*}$, the corresponding Chow functions for $P^{*}$
Theorem 2.16.
Suppose $P$ is a finite weakly ranked poset with a left KLS function $g$. Then
$A^{*}=\mathcal{I}(H^{T}),\ \ H^{*}=\mathcal{I}(A^{T})\ \ \mbox{ and }\ \ G^{*}=G^{T}.$
Proof.
By Theorem 2.11 for $P$,
$G=\mathcal{I}(g)H=Ag,\ \ \mathcal{I}(G)=G\ \ \mbox{ and }\ \ t\mathcal{I}(A)=A+(t-1)\delta.$
Transpose these identities and apply $\mathcal{I}$ to deduce
$G^{T}=\mathcal{I}(H^{T})g^{T}=\mathcal{I}(g^{T})\mathcal{I}(A^{T}),\ \ \mathcal{I}(G^{T})=G^{T}\ \ \mbox{ and }\ \ t\mathcal{I}(\mathcal{I}(H)^{T})=\mathcal{I}(H)^{T}+(t-1)\delta,$
where we have used Theorem 2.2 for the last equation. The theorem now follows from the uniqueness in Theorem 2.11 for $P^{*}$. ∎
@@ 4.2. Interlacing zeros and $\mathcal{I}_{n}$-interlacing sequences of polynomials
A polynomial $f\in\mathbb{C}[t]$ is called stable if either $f\equiv 0$ or all the zeros of $f$ have nonpositive imaginary parts. The Hermite-Biehler theorem relates stability to interlacing zeros, for a proof see [Wag11] [^4] There is a typo in the definition of proper position in [Wag11]. It is essential to add the condition that the zeros of $f$ and $g$ interlace..
@@ 5.2. Interlacing properties of $\gamma$-Chow polynomials
6. Chow polynomials of $\mathrm{TN}$-posets and dual $\mathrm{TN}$-posets
@@
    $q$-posets with nonnegative $h$-vectors [Ald10]. For example shellable $q$-complexes [GPR22].
@@
Suppose $P$ is a $\operatorname{\mathrm{TN}}$-poset, or the dual of a $\operatorname{\mathrm{TN}}$-poset, of rank $n$. Then the Chow polynomials $H_{P},d_{P},G_{P}$ and $A_{P}$ are real-rooted.
Moreover if $P$ is $\operatorname{\mathrm{TN}}$, then $H_{\tau(P)}\prec H_{P}$ and $H_{[\hat{0},x]}\prec H_{P}$ for any $x$ of rank n1n-1.
@@
Suppose $P$ is a $\operatorname{\mathrm{TN}}$-poset. Let $R=R(P)$. By Proposition 3.5, $H_{n}=H_{P}$, $d_{n}=tH_{\tau(P)}$ and $H_{n-1}=H_{[\hat{0},x]}$. Hence the statements for $\operatorname{\mathrm{TN}}$-posets follow from Theorems 4.18 and 4.19.
The case when $P^{*}$ is $\operatorname{\mathrm{TN}}$ follows from the above and Theorem 2.16, noting again that $d_{P}=tH_{\tau(P)}$ and that $\tau(P)$ is a dual $\operatorname{\mathrm{TN}}$-poset. ∎
If we apply Theorem 6.1 to a above, then we recover the result of Hoster and Stump [HS25]. Theorem 6.1 for the case e above implies in particular that Chow polynomials of Dowling lattices are real-rooted [^5] Coron, Ferroni, and Li informed us that they independently found a different proof for the real-rootedness of Chow polynomials of Dowling lattices..
@@
Let $P$ be a dual $\operatorname{\mathrm{TN}}$-poset. Then the characteristic Chow polynomial interlaces the characteristic augmented Chow polynomial, i.e., $H_{P}\prec G_{P}$.
Proof.
Let $P^{*}$ denote the poset dual to $P$. By Theorems 4.19 and 2.16 we deduce
$G_{P}=G_{P^{*}}\prec A_{P^{*}}=tH_{P},$
from which the result follows. ∎
@@ 8.1. Binomial and Sheffer posets
$\sum_{n=0}^{\infty}G_{n}(t)\frac{z^{n}}{C(n)}$ $=\frac{c(z)\cdot b(tz)-tc(tz)\cdot b(z)}{b(tz)-tb(z)}.$